OpenStudy (anonymous):

Please, PLEASE HELP! I really need it! Algebra 2! If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to y from the first equation is substituted into the second equation. 12x - y = -4 4x - 3y = -6 Answer choices: 4x - 3(-12x - 4) = -6 4x - 3(12x + 4) = -6 4(12x + 4) - 3y = -6 4(-12x - 4) - 3y = -6

OpenStudy (anonymous):

@tHe_FiZiCx99 Think you might be able to help? :l

OpenStudy (the_fizicx99):

Ohhh these, yes :P

OpenStudy (anonymous):

Awesome!

OpenStudy (the_fizicx99):

The substitution method is when you convert one of the equations to isolate y or x to then plug in, I'll find y, 12x - y = -4 -12x -12x -y = -12x -4 /-1 /-1 /-1 y = 12x + 4 Sub in y = 12x + 4 for y into the other equation, 4x - 3(12x + 4) = -6 Apply distributive property, 4x -36x -12 = -6 -32x -12 = -6 +12 +12 -32x = 6 /-32 /-32 x = -0.1875 OR -3/16 Plug in x to solve for y, 4(-0.1875) - 3y = -6 -0.75 - 3y = -6 +.75 +.75 -3y = -5.25 /-3 /-3 y = 1.75 OR 7/4

OpenStudy (mathmale):

AdrianK: Please start by solving 12x - y = -4 for y.

OpenStudy (the_fizicx99):

(-0.1875,1.75)

OpenStudy (anonymous):

That's not an answer choice though

OpenStudy (mathmale):

AdrianK: Please start by solving 12x - y = -4 for y. Once you have this expression for y, substitute it into the 2nd equation; let this expression take the place of y in the second expression.

OpenStudy (anonymous):

y=12x-4

OpenStudy (the_fizicx99):

Yes, it is. The question to solve for y and then plug in that for y into the other equation.

OpenStudy (the_fizicx99):

Whatever y equal's after you solved for y, plug that in, plug in " y=12x-4" into 4x - 3y = -6 for y 4x - 3( ?? ) = -6 Get it?

OpenStudy (the_fizicx99):

It's in my illustration above lol

OpenStudy (anonymous):

I kinda get it. It just dont see your solutions though

OpenStudy (anonymous):

like 1.75 or 7/4 isn't an answer

OpenStudy (the_fizicx99):

What do you mean solution? I completely solved it. Though they didn't ask me to, they only asked " which of these equation looks like after plugging in for y." Meaning after you substitute your y into the other equation, pick which is the one that looks exactly like it. Suppose they asked for: 3x + 2y = 2 and 5x + 8y = 2 I would solve for y, Take the first equation and solve for y, y = -3/2x + 1 Plug that into Y for 5x + 8y = 2 It looks like this: 5x + 8(-3/2x + 1) = 2

OpenStudy (anonymous):

so it would be 2 and -3?

OpenStudy (the_fizicx99):

Look at your options, after you solved the first equation for y, plug that in for 4x - 3y= -6

OpenStudy (anonymous):

OH MY GOD! I'm sorry!!

OpenStudy (anonymous):

I was looking at the wrong question! I was looking at the oone above this one -_-

OpenStudy (anonymous):

so it would be B?

OpenStudy (the_fizicx99):

Just substitute the y you solve for, into 4x-3(SUB the Y you solved for) = -6

OpenStudy (the_fizicx99):

Yes, B :P

OpenStudy (anonymous):

again, im sorry lol

OpenStudy (the_fizicx99):

np, its fine :P

OpenStudy (anonymous):

Think you could help me with one more? (:

OpenStudy (the_fizicx99):

ya

OpenStudy (anonymous):

What are the x-coordinates of the solutions to this system of equations? x2 + y2 = 16 y = x + 4 Answer Choices: -4 and 0 4 and 0 -2 and -3 2 and -3

OpenStudy (the_fizicx99):

-4 and 0

OpenStudy (the_fizicx99):

to solve for x set y for 0 to solve for y set x for 0

OpenStudy (anonymous):

Thanks!!!

OpenStudy (the_fizicx99):

np